A chemist has 50ml of a solution with 20% concentration of acid. How many ml of pure acid must be added to increase the concentration to 30%?
Answer by richwmiller(17219)

1*x+0.2*50=0.3(50+x)
1*x+10=15+0.3x
1x-0.3x=15-10
0.7x=5
x=7.14285714
check
1*7.14285714+0.2*50=0.3(50+7.14285714)
7.14285714+10=0.3(57.1428571)
17.1428571=17.1428571
ok